OpenStudy (anonymous):

Can someone please explaine ratios and solving proportions? Thanksss(:

OpenStudy (anonymous):

Ratios and proportions are fractions. Remember equivalent fractions? 1/2 =2/4=5/10=14/28 Well ratios are fractions, and since they're usually maintained, you can compare different values as long as you keep the same ratio. Do you have a specific problem to look at?

OpenStudy (anonymous):

um, 18/20 K/110... it says to use cross products! What is that?

OpenStudy (anonymous):

@Rudy

OpenStudy (anonymous):

So your ratio is 18/20 and since a ratio is maintained, you can say that it equals k/110 so: \[\frac{18}{20}=\frac{k}{110}\] Which you can solve as any equation. To isolate k, multiply times 110 on both sides. \[\frac{18*110}{20}=\frac{k*110}{110}\] \[\frac{18*110}{20}=k\]

OpenStudy (anonymous):

there are others but i woll try my best on them! I just dont under stand this one!

OpenStudy (anonymous):

*Will

OpenStudy (anonymous):

so its 1,800?

OpenStudy (anonymous):

haha, thats way off!

OpenStudy (anonymous):

Just remember a ratio is maintained. So you can equate the ratios given, no matter the change in the amount... Like unit prices. If someone says a quart of oil is 3.25 how many can you buy with 22.75 Well the fraction is: \[\frac{quarts of oil}{price}=\frac{1}{3.25}=\frac{x - quarts}{22.75}\] Hope that helps. Let me know if you have any you're stuck on, I may take time to answer, but I will.

OpenStudy (anonymous):

oh!! So thats the cross thing? Not that hard, thank you so much!! All of them are alike!! haha, thanks so much!!

OpenStudy (anonymous):

you're welcome.